OpenStudy (anonymous):

you invest a total of $13750 in two companies (part in SLW and part in STDY). After one year the SLW investment increases by 6.4% and the STDY investment increases by 3.9%. How much was invested in STDY if the TOTAL increase in the initial investment was 6.2%

OpenStudy (amistre64):

so if the total increase of the initial investment (from both companies? or just from this one) is 6.2%; then ....

OpenStudy (anonymous):

so 6.2 =....

OpenStudy (anonymous):

am i right so far?

OpenStudy (amistre64):

x + y = 13750 x(6.4%) + y(3.9%) = 13750(6.2%) we good with this so far?

OpenStudy (anonymous):

yea and i can take it from there

OpenStudy (anonymous):

i can just use the elimination method right?

OpenStudy (amistre64):

id rename it SLW and STDY just to be on the safe side :)

OpenStudy (amistre64):

elimination is when you canel out one variable...it might be an option; but usually I just solve for one of the top variables and sustitute it into the bottom one.

OpenStudy (anonymous):

that would be easier :)

OpenStudy (amistre64):

x + y = 13750 y = 13750 - x --------------------------- x(.064) + y(.039) = 13750(.062) x(.064) + (13750 - x)(.039) = 13750(.062)

OpenStudy (amistre64):

solve for "x" and then back it up for "y"

OpenStudy (anonymous):

right. i will tell you what i get, one minute

OpenStudy (anonymous):

i got 1.59 for x now about to plug back in for y

OpenStudy (anonymous):

i think i did something wrong thats REALLY small

OpenStudy (amistre64):

ok....ill double check your work :)

OpenStudy (anonymous):

im lost

OpenStudy (amistre64):

x = 12650

OpenStudy (anonymous):

i multiplied everything

OpenStudy (anonymous):

and got .064x + 536.25x = 852.5

OpenStudy (anonymous):

solved for x and got 1.59 (rounded) but that doesnt work when i solve for y

OpenStudy (amistre64):

x(.064) + (13750 - x)(.039) = 13750(.062) .064x -.039x + 536.25 = 852.50 .025x = 316.25 x = 316.25/.025 x = 12650

OpenStudy (anonymous):

OHH i didn't distribute the .039 properly

OpenStudy (anonymous):

can you just clarify for me and show me kinda step by step how you set up the two equations? based off of the word problem?

OpenStudy (amistre64):

I can..if your still around

OpenStudy (anonymous):

yes

OpenStudy (amistre64):

There are 2 companies that you invest a certain number of dollars in: one company get s "x" amount of dollars and the other gets "y" amount of dollars. the amount invested into both companies has to equal 13750. x + y = 13750

OpenStudy (anonymous):

right

OpenStudy (amistre64):

We know that each company has a different interest rate that they pay so we set this up accordingly we know that x amount of dollars gets us 6.4% and y amount of dollars gets us 3.9%. We are also told that the total amount of interest earned from all the monies invested was 6.2%.

OpenStudy (amistre64):

interest earned is a simple calcultion; it is amount of money times interest rate. plain and simple. x(6.4%) + y(3.9%) = 13750(6.2%) right?

OpenStudy (anonymous):

yes, ok i get it i just wanted to reinforce it

OpenStudy (amistre64):

ok :)